OpenStudy (anonymous):

Teri and Jennifer have 480 holiday cookies altogether. Teri has 1 and 2/3 as many cookies as Jennifer. How many cookies does Jennifer have?

OpenStudy (bemoseemo):

is there a list of answers or is it one of those that you have to guess? ¯\_(ツ)_/¯

OpenStudy (anonymous):

Nope :(

OpenStudy (anonymous):

do you have anyone helping you?

Vocaloid (vocaloid):

let's define our variables first let T = the number of cookies terri has let J = the number of cookies jennifer has with me so far?

OpenStudy (anonymous):

i am not sure if he is still here, he has not responded to me at all

OpenStudy (anonymous):

I'm here sorry

OpenStudy (anonymous):

no problem, but i think that the expert left..

OpenStudy (anonymous):

my computer doesn't update fast when someone responds all the time

OpenStudy (anonymous):

that stinks

Vocaloid (vocaloid):

I'm back lol ok, so they have 480 cookies altogether, then T + J = 480, with me so far?

OpenStudy (anonymous):

yes!

Vocaloid (vocaloid):

it also says that terri has 1 2/3 as many cookies as jennifer to make things easier, I'm going to rewrite 1 2/3 as 5/3, since these two fractions are equal so T = (5/3)J, with me so far?

OpenStudy (anonymous):

yes!

Vocaloid (vocaloid):

great, now we can replace "T" with "5/3 J" so T + J = 480 becomes J + (5/3)J = 480 know where to go from here?

OpenStudy (anonymous):

would you do 2j x 5/3= whatever

Vocaloid (vocaloid):

where are you getting 2 from

Vocaloid (vocaloid):

start by adding J + (5/3)J = ?

OpenStudy (anonymous):

combining the 2 js

Vocaloid (vocaloid):

uh not quite 1 + (5/3) = ?

OpenStudy (anonymous):

2.66666667

Vocaloid (vocaloid):

as a fraction please

OpenStudy (anonymous):

terri and jennifer now have diabetes

OpenStudy (anonymous):

2 2/3

Vocaloid (vocaloid):

good, we can also write 2 2/3 as 8/3 now we have J + (5/3)J = (8/3)J so (8/3)J = 480 know where to go from here?

OpenStudy (anonymous):

@Vocaloid i feel as if you are a genius

OpenStudy (anonymous):

no kidding

OpenStudy (anonymous):

divide 8/3 on both sides

Vocaloid (vocaloid):

yup so 480/(8/3) = ?

OpenStudy (anonymous):

20?

Vocaloid (vocaloid):

hmmm not quite when we divide by a fraction, we can flip the fraction and multiply so 480/(8/3) becomes 480*(3/8) which equals...?

OpenStudy (anonymous):

180

Vocaloid (vocaloid):

ahaha that's nice of you guys to say, I'm far from a genius though

Vocaloid (vocaloid):

and yes 180 is the answer, good job~

OpenStudy (anonymous):

thank you so much!

Vocaloid (vocaloid):

so Jennifer has 180 cookies

OpenStudy (anonymous):

wait... if you have time, can you help me?